# Principal-Agent Multitasking: the Uniformity of Optimal Contracts and its Efficient Learning via Instrumental Regression

### v1

@@ 3 Multitask Principal-Agent: Problem Description and the Uniformity Result
A function $c$ is homogeneous of degree $k$, if for any $\rho>0$ and $w\in(\mathbb{R}^{d})^{+}$, $c(\rho w)=\rho^{k}c(w)$.
@@
Definition 2.
The cost function $c$ is said to satisfy the Inada conditions if the following holds.
  1. 1.
    $c$ is convex on its domain.
  2. 2.
    $\lim_{w_{i}\rightarrow 0}\partial c(w)/\partial w_{i}=0$.
  3. 3.
    $\lim_{w_{i}\rightarrow+\infty}\partial c(w)/\partial w_{i}=+\infty$.
The Inada condition is fairly standard. It essentially means the marginal cost goes from 00 to +$\infty$+ ∞ as $w_{i}$ goes from 00 to +$\infty$+ ∞. Throughout the following, the following assumption will be imposed on the cost function of the agent(s).
@@
The cost function of the agent is homogeneous of degree $k$, and satisfies the Inada conditions.
@@ 3.1 The Uniformity Result
This subsection shows the optimal contract depends on the shape of $c(\cdot)$ only through the degree of homogeneity.
@@
The principal’s optimal contract is $\beta=\theta/k$.
@@
When the posted contract is $\beta$, the agent’s expected utility when taking action $w$ is
$\langle\beta,w\rangle-c(w).$
The first-order condition gives:
$\beta=\nabla c(w).$
The first-order condition must bind under 1.
The principal’s expected utility, written as a function of $w$, is then
$\langle\theta,w\rangle-\langle\beta,w\rangle=\langle\theta,w\rangle-\langle\nabla c(w),w\rangle=\langle\theta,w\rangle-kc(w).$
Note the above is a concave function in $w$. The first-order condition gives
$\theta=k\nabla c(w).$
But also recall at the agent’s optimum $\beta=\nabla c(w)$. Hence $\beta=\theta/k$ is the contract that maximizes the principal’s utility. ∎
@@ 3.2 Discussion of the Result
As a special case, the results explain the widespread 50-50 split rule seen in many scenarios (e.g., [22]). Specifically, when the agents have a quadratic cost function ($k=2$), the results indicate that the 50-50 split is precisely the optimal linear contract which ensures maximum utility for the principal.
4 Estimating the Optimal Contract: an Instrumental Regression Approach
@@
[…]
@@
     // If using small random perturbation for ‘exploration’:
     Set $\beta_{t}=(\theta_{0}+\rho_{t})/k$, here $\rho_{t}$ is a suitable random vector with variance $\sigma_{0}^{2}$ in each coordinate
@@
I now study a setting where the parameter $\theta$ is unknown, i.e., the marginal utility of each task is unknown. Stated a bit informally, the principal is unsure about the exact “importance” of each task. For example, when the manager (the agent) manages a firm and participates in multiple tasks, the firm’s owner (the principal) may be unsure of the exact importance of each task and how much each task contributes to the firm’s total revenue. When the parameter $\theta$ is unknown, the principal must estimate or learn this parameter through observed data. By the previous result (Theorem 1), which states the optimal contract is $\theta/k$, finding an approximately optimal contract is essentially equivalent to obtaining an accurate estimate for $\theta$. To make matters a bit simpler, assume the return to scale parameter $k$ is known and treated as a constant.
The data consists of tuples of the form $(\beta_{t},w_{t},x_{t},y_{t})$, representing the contract, agent’s true effort, observed signal, and total production outcome respectively. In the following, the total production outcome $y_{t}$ will just be referred to as the total revenue for short. Note that the observed data by the principal only include $(\beta_{t},x_{t},y_{t})$, and the agent’s true effort $w_{t}$ is hidden.
$\mathbb{E}[x_{t}|w_{t}]=w_{t}$
and that the total revenue satisfies a linear relation:
$\mathbb{E}[y_{t}|w_{t}]=\langle\theta,w_{t}\rangle.$
Fix any agent’s type (i.e., a cost function that satisfies 1). Define $u(\beta)$ as the expected utility when posting contract $\beta$. The principal is interested in finding a contract $\beta$ that minimizes the utility loss, which is the expected utility difference between posting the optimal contract $\theta/k$ and the contract $\beta$:
$u(\theta/k)-u(\beta).$
The following two assumptions are imposed, which allows us to relate the utility loss to the squared estimation error.
Assumption 2.
$c_{0}\leq\theta_{i}\leq C_{0}$ for $i\in[d]$.
The above assumption is natural and relatively mild, it states that there is some lower and upper-bound on the marginal utility of each task.
Assumption 3.
There exists some constant $c$ so that for any $\hat{\theta}\in[c_{0},C_{0}]^{n}$, the following holds
|$\absolutevalue{u(\theta/k)-u(\hat{\theta}/k)}\leq c\cdot\norm{\theta/k-\hat{\theta}/k}_{2}^{2}.$
The above assumption can essentially be seen as a smoothness assumption on the function $u$. Then, if the principal has an estimate $\hat{\theta}$ and the principal posts $\beta=\hat{\theta}/k$ as the contract, the utility loss can be bounded by the square estimation error (up to constant factors):
|$\absolutevalue{u(\theta/k)-u(\hat{\theta}/k)}\leq O\left(\norm{\theta-\hat{\theta}}_{2}^{2}\right).$
The above will hold for any agent type as long as the assumptions are met. Therefore, the following will mainly focus on obtaining upper bounds on the quantity $\norm{\theta-\hat{\theta}}_{2}^{2}$.
The problem of finding the optimal contract can either take place in an offline or online setting. In the offline setting, a sequence of data consisting of observations $(\beta_{t},x_{t},y_{t})$ is given to the principal, and the principal must obtain an accurate estimate of the true parameter $\theta$.
In the online setting, the principal interacts with agents in a sequential fashion consisting of $T$ rounds. In this scenario, for each round $t$, the principal must choose the contract $\beta_{t}$ based on historical data, then observe the signals $x_{t}$ and the total revenue $y_{t}$. In this online setting, the principal may be interested in minimizing the cumulative utility loss. A difference from the offline setting is that the principal needs to engage in some ‘experimentation’. Specifically, the principal may want to strategically choose some contracts $\beta$ in order to get a better estimate of the true parameter $\theta$; however, this may come at the cost of losing immediate revenue.
4.1 Offline Setting: Estimation with Instrumental Regression
The causal relationship between $(\beta_{t},w_{t},x_{t},y_{t})$ is summarized in Figure 1. The problem of estimating $\theta_{t}$ is essentially a (linear) regression problem. It can be seen that the dependence between the variables form a measurement error model, and it is well-known that directly regressing on $x_{t}$ as an attempt to learn $\theta$ will lead to a biased estimate. However, it can be observed that the contract $\beta_{t}$ acts as a valid instrumental variable. Specifically, it can be observed that
$\mathbb{E}[\beta_{t}(y_{t}-\langle\theta,x_{t}\rangle)]=\mathbf{0}\ (\in\mathbb{R}^{d}).$
$\hat{\theta}_{T}=(B^{\top}_{T}X_{T})^{-1}B^{\top}_{T}Y_{T}.$
Here $B_{T}\in\mathbb{R}^{T\times d}$ is a matrix with rows representing the posted contracts each round; $X_{T}\in\mathbb{R}^{T\times d}$ is a matrix with rows representing the signals each round; $Y_{T}\in\mathbb{R}^{T}$ is a column vector representing the outcomes each round. Denote $\gamma_{t}=y_{t}-\langle\theta,x_{t}\rangle$, and collect values for $\gamma_{1:t-1}$ into the column vector $\Gamma_{t}$.
@@ 4.1 Offline Setting: Estimation with Instrumental Regression
The estimation error
$\norm{\theta-\hat{\theta}_{T}}_{2}\leq\frac{\norm{B_{T}^{\top}\Gamma_{T}}}{\sigma_{\min}(B_{T}^{\top}X_{T})}.$
Proof.
$\hat{\theta}_{T}$ $=(B_{T}^{\top}X_{T})^{-1}B_{T}^{\top}Y_{T}$
$=(B_{T}^{\top}X_{T})^{-1}B_{T}^{\top}(X_{T}\theta+\Gamma_{T})$
$=\theta+(B_{T}^{\top}X_{T})^{-1}B_{T}^{\top}\Gamma_{T}.$
The proposition then follows. ∎
@@
By a standard concentration inequality, the numerator can be upper bounded as $O(\sqrt{d{T}})$ with high probability. The estimation error then depends crucially on the minimum singular value of $B_{T}^{\top}X_{T}$.
4.2 Online Setting: Learning the Optimal Contract
The estimator above gives a estimation error bound when offline observations are available. However, it does not directly give rise to a learning algorithm in the online setting, where the principal needs to choose the contract each round. To give some intuition, in this setting, the principal essentially faces an optimal design problem with an exploration-exploitation tradeoff. Specifically, the matrix $B_{t}$, consisting of contracts $\beta_{t}$, is the design, and the principal must choose this matrix such that the estimate $\hat{\theta}$ becomes accurate. At the same time, the principal must ensure that the chosen contracts $\beta_{t}$ are close to the true optimal contract $\theta/k$ in order to guarantee a small cumulative utility loss.
Lower bound on regret: $\Omega(d\sqrt{T})$
Consider a setting where each agent has cost function $c(w)=\frac{1}{2}\norm{w}_{2}^{2}$, and that this cost function is known to the principal. Then, the agent’s best response under the contract $\beta$ is exactly $w=\beta$. The principal’s expected utility when a contract $\beta$ is posted is then
$u(\beta)=\langle\theta,\beta\rangle-\norm{\beta}_{2}^{2}.$
The optimal contract is then $\beta=\theta/2$ (this is also implied by the previous uniformity result). The principal now essentially faces a zero-order optimization problem with a quadratic objective function. The contract $\beta$ is the decision variable, and the realized total revenue is the observed variable. Further, by a result in [23], the tight regret bound for this problem is $\Omega(d\sqrt{T})$.
Simple algorithms achieving $O(d\sqrt{T})$ cumulative utility loss
Perhaps a bit surprisingly, a very simple algorithm can achieve $O(d\sqrt{T})$ regret. In the first $d\sqrt{T}$ rounds, for each $i\in[d]$, the principal should set the contract as $\beta=\mathbf{e}_{i}$ for $\sqrt{T}$ rounds. The agents best response $w\in\mathbb{R}^{d}$ should only have a non-zero value in the $i$-th component $w_{i}$. Hence, the data in these rounds can be used to estimate $\theta_{i}$, and the estimation error after the initial $d\sqrt{T}$ rounds is:
|$\absolutevalue{\hat{\theta}_{i}-\theta_{i}}=O(1/{T^{1/4}}).$
In the remaining rounds, the principal can simply use the estimate after the initial $d\sqrt{T}$ rounds, and post the contract as $\beta=\hat{\theta}/k$. The regret in these rounds can be bounded as:
$T\cdot\norm{\hat{\theta}-\theta}_{2}^{2}=d\sqrt{T}.$
@@ 4.2 Online Setting: Learning the Optimal Contract
5 Improving the Estimation with Repeated Observations and Diversity
@@
Epoch $e$ contains |$\absolutevalue{\tau_{e}}=\max\{d\cdot 2^{e},8dC_{0}K_{0}\ln(d/\delta)/(c_{0}^{2}\lambda_{0}),4\sigma_{0}^{2}d^{2}\ln(d^{2}/\delta)/(c_{0}^{2}\lambda_{0})\}$ rounds
@@
This section continues the discussion where the principal must learn $\theta$ using sequential observations. The previous section showed that using an instrumental regression approach where the contract is the instrumental variable, the principal must balance exploration (ensuring that $\sigma_{\min}(B_{t}^{\top}X_{t})$ grows as $t$ increases) and exploitation (ensuring $\beta_{t}$ is close to the true optimal contract). In this section, I show that, under some additional assumptions, the learning rate of the principal can be greatly improved. Informally, I show that when the agents are sufficiently ‘diverse’, and that when repeated observations are available, the principal can achieve a logarithmic regret, even if the principal is using a ‘pure exploitation algorithm’.
To state the results in a more concise manner, this section restricts attention to the case where the cost functions of the agents have diagonal quadratic forms. Specifically, the cost function of the agent is parameterized by a vector $\kappa\in(\mathbb{R}^{d})^{+}$, so that the cost function of the agent is $c(w)=\frac{1}{2}w^{\top}\mathtt{diag}(\kappa)^{-1}w$. Then, the agent’s best response to a contract $\beta$ is the maximizer of $\langle\beta,w\rangle-c(w)$, which is equal to
$w=\kappa\odot\beta.$
I assume that the agent cost function in each round, parameterized by $\kappa_{t}$, is drawn from an unknown distribution. The distribution satisfies the following.
@@
The minimum eigenvalue of $\mathbb{E}_{\kappa\sim\mathcal{D}}[\kappa\kappa^{\top}]$ is lower bounded by $\lambda_{0}$. Further, for any $\kappa$ in the support of $\mathcal{D}$, $\kappa\leq K_{0}$.
The key part in the above assumption is the minimum eigenvalue $\lambda_{0}$, it suggests that the agents’ are sufficiently “diverse" in their types. The second assumption is that repeated observations are available. I assume that for each agent, the principal can observe two sets of signals, denoted by $x_{t}$ and $\widetilde{x}_{t}$. The signals are $\sigma_{0}^{2}$-subgaussian vectors and satisfy the following:
$\mathbb{E}[x_{t}|w_{t}]=\mathbb{E}[\widetilde{x}_{t}|w_{t}]=w_{t}.$
@@
The two sets of signals are assumed to be (conditionally) independent. Specifically:
$x_{t}\perp\widetilde{x}_{t}|w_{t}.$
The causal relationship is summarized in Figure 1 with the dashed items included. It can be observed that the second set of observations can serve as an instrumental variable for the other, specifically:

### v2

@@ 5 Uniformity of the Optimal Linear Contract
The cost function of the agent is homogeneous of degree $k$. In other words, for any $\rho>0$ and $a\in(\mathbb{R}^{d})^{+}$, $c(\rho a)=\rho^{k}c(a)$.
Remark 2.
@@
We also assume the private benefit to the principal depends only on the agent’s true effort vector $a$ through a linear model.
@@
There exists $\theta^{*}\in(\mathbb{R}^{d})^{+}$, such that
$\mathbb{E}[y(a)]=\langle\theta^{*},a\rangle.$
The value $\theta^{*}_{i}$ can be interpreted as the marginal utility of task $i$ to the principal.
@@ 5.1 The Uniformity Result
Here we show the optimal linear contract depends on the shape of $c(\cdot)$ only through the degree of homogeneity. The proof is extremely simple, yet to the best of our knowledge, it has not appeared in prior work.
@@
The principal’s optimal contract is $\beta^{*}=\theta^{*}/k$.
@@
Instead of finding the best optimal linear contract, let us find the best action of the agent to induce. Suppose the best action to induce were $a$. By Lemma 1, the best linear contract to induce $a$ is the contract $\beta=\nabla c(a)$. In other words, $\beta=\nabla c(a)$ induces $a$, and no other contracts induces $a$ at a lower payment. Then, the principal’s expected utility for inducing action $a$ is
$\langle\theta^{*},a\rangle-\langle\beta,a\rangle$ $=\langle\theta^{*},a\rangle-\langle\nabla c(a),a\rangle$
$=\langle\theta^{*},a\rangle-kc(a).$
The above is a concave function of $a$, and the first-order condition gives:
$\theta^{*}=k\nabla c(a).$
But also recall we have $\beta=\nabla c(a)$. Hence at the optimal we must have $\beta^{*}=\theta^{*}/k.$
@@ 5.2 Discussion of the Result
A problem setting very similar to the current work was studied in the recent work [Roth et al., 2016]. Their work studied optimization algorithms in Stackelberg games from revealed preferences. As a special case, they gave a discussion on how their algorithm can be applied to the multitask principal-agent problem (though their work did not explicitly identify their problem as such). However, the uniformity result which we presented here seemed to have somehow escaped their analysis.
6 Estimating and Learning the Optimal Contract: an Instrumental Regression Approach
@@
This section builds upon the results in the previous section where we presented the uniformity result, but consider a setting where the parameter $\theta^{*}$ is unknown so that some estimation or learning is involved to obtain the optimal linear contract. In other words, the principal lacks precise knowledge of the relative ”importance” of each task. For instance, when a firm’s owner (the principal) hires a manager (the agent) to oversee multiple tasks, the owner may be uncertain about the exact contribution of each task to the firm’s overall performance. In such cases, the principal must estimate or learn $\theta^{*}$ using observational data. By the uniformity result (Theorem 2), which establishes that the optimal contract is given by $\beta^{*}=\theta^{*}/k$, identifying an approximately optimal contract reduces to obtaining an accurate estimate of $\theta^{*}$.
We assume that the homogeneity degree parameter $k$ is known and treated as a constant. In other words, the return to scale parameter is a constant for different agent types. While one could also study the scenario where $\theta^{*}$ is known but $k$ is unknown, this case is less technically interesting. Since $k$ is just a single-dimensional parameter, it can be inferred using a some one-dimensional discretization-based algorithm. Thus, we focus on the case where $\theta^{*}$ is unknown while $k$ is known and fixed, which seems to be the more technically interesting problem.
Note that while this bears some resemblance to the linear bandit problem, where seemingly $x$ represents the context and the private benefit $y$ represents the noisy realized reward, our problem is in fact quite different. This is because, in fact, the signal is a noisy realization of the true “context”, which is the agent’s private effort. Recall that $\mathbb{E}[y(a)]=\langle\theta^{*},a\rangle$, so $a$ is the true “context” vector in the linear bandit problem. Essentially, our problem faces a measurement error problem.
We assume the principal has access to data in $T$ time periods. In time period $t$, the observed data consists of tuples of the form $(\beta_{t},x_{t},y_{t})$, representing the contract, observed signal, and private benefit respectively; the agent’s private effort $a_{t}$ is unobserved by the principal. The relationship is depicted in Figure 1.
$\mathbb{E}[x_{t}|a_{t}]=a_{t},$
and that the total revenue satisfies a linear relation:
$\mathbb{E}[y_{t}(a_{t})]=\langle\theta^{*},a_{t}\rangle.$
We make a mild assumption that coordinates of $\theta^{*}$ are bounded, so that there is some bounded set $\mathcal{B}$ in which the optimal contract $\beta^{*}$ belongs to. This is relatively mild. Even though the principal may not know the exact value of $\theta^{*}$, she may have some reasonable bound on the marginal utility of each task, i.e., on $\theta^{*}_{i}$. For example, if the principal knows that each $\theta_{i}$ will be bounded in the interval $[c_{0},C_{0}]$, then we can take $\mathcal{B}=[c_{0}/k,C_{0}/k]^{d}$. Without loss of generality we can perform normalization and assume the following holds.
Assumption 3.
There exists some set $\mathcal{B}\subset[0,1]^{d}$, such that $\beta^{*}\in\mathcal{B}$.
Moreover, we assume the signals $x_{t}$ and realized private benefits $y_{t}$ are conditionally subgaussian.
Assumption 4.
Given the agent’s action $a_{t}$, the random signal $x_{t}$ and private benefit $y_{t}$ are $\sigma$-subgaussian.
The problem of determining the optimal contract can be approached in either an offline or an online setting. In the offline setting, the principal is given a sequence of data consisting of observations $(\beta_{t},x_{t},y_{t})$ and must estimate the true parameter $\theta^{*}$ as accurately as possible.
In the online setting, the principal interacts with agents sequentially over $T$ rounds. In each round $t$, the principal selects a contract $\beta_{t}$ based on past observations and then observes the corresponding signals $x_{t}$ and total revenue $y_{t}$. Unlike in the offline case, the principal now face a exploration vs exploitation tradeoff: while better estimates of $\theta^{*}$ improve contract effectiveness in the long run, acquiring these estimates often requires experimentation, which may lead to short-term revenue loss.
6.1 Offline Setting: GMM Estimator
The causal relationship between $(\beta_{t},a_{t},x_{t},y_{t})$ is summarized in Figure 1. The problem of estimating $\theta_{t}$ is essentially a (linear) regression problem. It can be seen that the dependence between the variables form a measurement error model, and it is well-known that directly regressing on $x_{t}$ as an attempt to learn $\theta^{*}$ will lead to a biased estimate. However, it can be observed that the contract $\beta_{t}$ acts as a valid instrumental variable. Specifically, it can be observed that the following moment condition is satisfied
$\mathbb{E}[\beta_{t}(y_{t}-\langle\theta^{*},x_{t}\rangle)]=\mathbf{0}\ (\in\mathbb{R}^{d}).$
$\hat{\theta}_{T}=(B^{\top}_{T}X_{T})^{-1}B^{\top}_{T}Y_{T}.$
Here $B_{T}\in\mathbb{R}^{T\times d}$ is a matrix with rows representing the posted contracts each round; $X_{T}\in\mathbb{R}^{T\times d}$ is a matrix with rows representing the signals each round; $Y_{T}\in\mathbb{R}^{T}$ is a column vector representing the outcomes each round.
Proposition 1.
With probability at least $1-\delta$, the estimation error
$\norm{\theta^{*}-\hat{\theta}_{T}}_{2}\leq\frac{\sqrt{dT\log(dT/\delta)}}{\sigma_{\min}(B_{T}^{\top}X_{T})}.$
Proof.
Denote $\gamma_{t}=y_{t}-\langle\theta^{*},x_{t}\rangle$, and collect values for $\gamma_{1:t-1}$ into the column vector $\Gamma_{t}$.
$\hat{\theta}_{T}$ $=(B_{T}^{\top}X_{T})^{-1}B_{T}^{\top}Y_{T}$
$=(B_{T}^{\top}X_{T})^{-1}B_{T}^{\top}(X_{T}\theta^{*}+\Gamma_{T})$
$=\theta^{*}+(B_{T}^{\top}X_{T})^{-1}B_{T}^{\top}\Gamma_{T}.$
Hence
$\norm{\theta^{*}-\hat{\theta}_{T}}_{2}\leq\frac{\norm{B_{T}^{\top}\Gamma_{T}}}{\sigma_{\min}(B_{T}^{\top}X_{T})}.$
By a standard concentration inequality (see Section A.1) the numerator can be upper bounded as $\sqrt{dT\log(dT/\delta)}$ with probability at least $1-\delta$, the proposition then follows. ∎
6.2 Online Setting: Explore then Commit Learning Algorithm
@@
Number of exploration rounds $\tau=d\sqrt{T}$, an initial estimate $\theta_{0}$ for $\theta$
@@
In the previous part, we developed an estimator based on GMM for estimating the parameter $\theta^{*}$, and by extension (in particular by the uniformity result), the optimal contract $\beta^{*}$. This applies to the setting when offline observations are available. However, it does not directly give rise to a learning algorithm in the online setting, where the principal needs to choose the contract each round. To give some intuition, in this setting, the principal essentially faces an optimal design problem with an exploration-exploitation tradeoff. Specifically, the matrix $B_{t}$, consisting of contracts $\beta_{t}$, is the design, and the principal must choose this matrix such that the estimate $\hat{\theta}$ becomes accurate. At the same time, the principal must ensure that the chosen contracts $\beta_{t}$ are close to the true optimal contract $\beta^{*}=\theta^{*}/k$ in order to guarantee a small cumulative utility loss. We will be interested in bounding the following quantity, which is the square error of $\beta_{t}$ each round:
$\mathtt{Reg}=\sum_{t=1}^{T}\norm{\beta^{*}-\beta_{t}}_{2}^{2}.$
The above can serve as a proxy for the actual cumulative utility loss: $\sum_{t=1}^{T}\absolutevalue{u_{t}(\beta^{*})-u_{t}(\beta_{t})}$ |. In particular, as long as the function $u_{t}$ has a bounded second derivative at the global maximum $\beta^{*}$, the cumulative utility loss can be upper bounded by $O(\tt Reg)$.
6.2.1 Upper Bounds
Here we show that very simple explore-then-commit type algorithms can achieve $O(d\sqrt{T})$ regret. In the first $d\sqrt{T}$ rounds, for each $i\in[d]$, the principal should set the contract as $\beta=\mathbf{e}_{i}$ for $\sqrt{T}$ rounds. The agents best response $a\in\mathbb{R}^{d}$ should only have a non-zero value in the $i$-th component $a_{i}$. Hence, the data in these rounds can be used to estimate $\theta_{i}$. In the remaining rounds, the principal can simply use the estimate $\hat{\theta}$ after the initial $d\sqrt{T}$ rounds, and post the contract as $\beta=\hat{\theta}/k$.
@@ 6.2.1 Upper Bounds
There exists algorithm achieving regret $\widetilde{O}(d\sqrt{T})$ with probability $1-1/T$.
The proof of the above can be found in Section A.2.
@@
6.2.2 Lower Bound: $\Omega(d\sqrt{T})$
Consider a setting where each agent has cost function $c(a)=\frac{1}{2}\norm{a}_{2}^{2}$, and that this cost function is known to the principal. Then, the agent’s best response under the contract $\beta$ is exactly $a=\beta$. The principal’s expected utility when a contract $\beta$ is posted is then
$u(\beta)=\langle\theta^{*},\beta\rangle-\norm{\beta}_{2}^{2}.$
The optimal contract is then $\beta^{*}=\theta^{*}/2$ (this is also implied by the previous uniformity result, since the agent’s cost function is quadratic and homogeneous of degree $2$). The principal now essentially faces a zero-order optimization problem with a quadratic objective function. The contract $\beta$ is the decision variable, and the realized total revenue is the observed variable. Further, by a result in [Shamir, 2013], the tight regret bound for this problem is $\Theta(d\sqrt{T})$.
Proposition 3.
For any (possibly randomized) strategy the principal can adopt for posting contracts, there exists some parameter $\theta^{*}$ such that the term $\tt Reg$ can be lower bounded by $\Omega(d\sqrt{T})$.
7 Improving Estimation and Learning with Repeated Observations and Diversity
@@
Epoch $e$ contains |$\absolutevalue{\tau_{e}}=\max\{d\cdot 2^{e},8K_{0}\ln(d/\delta)/(\lambda_{0}),4\sigma_{0}^{2}d^{2}\ln(d^{2}/\delta)/(\lambda_{0})\}$ rounds
@@
This section continues the discussion from the previous section, where the principal must learn or estimate $\theta^{*}$ using observational data. The previous section showed that using an instrumental regression approach and using the contract as the instrumental variable, the principal must balance exploration (ensuring that $\sigma_{\min}(B_{t}^{\top}X_{t})$ grows as $t$ increases) and exploitation (ensuring $\beta_{t}$ is close to the true optimal contract). In this section, we show that, under some additional assumptions, the learning rate of the principal can be greatly improved. Informally, we show that when the agents are sufficiently ‘diverse’, and that when repeated observations are available, the principal can achieve a logarithmic regret, even if the principal is using a ‘pure exploitation’ algorithm.
[…]
Maybe clash between $B_{t}$ and $B$.
In addition to the boundedness of the the possible contract space $B$ in the previous section, we make two extra critical assumptions in this section. The first is that repeated signals are available. The second is that agents are drawn from some stochastic distribution and are sufficiently diverse.
In the following, let us suppose that agents are drawn from some distribution $\mathcal{D}$, each element in the support of $D$ represents the cost function of the agent. It is assumed that for all elements in the support of $D$, the cost function is homogeneous of degree $k$ (i.e., it satisfies 1). Our first part of the assumption below states that agents are sufficiently diverse.
@@
Repeated observations are available. We assume that in each round, the principal can observe two sets of signals, denoted by $x_{t}$ and $\widetilde{x}_{t}$. The signals are $\sigma_{0}^{2}$-subgaussian vectors and satisfy the following:
$\mathbb{E}[x_{t}|a_{t}]=\mathbb{E}[\widetilde{x}_{t}|a_{t}]=a_{t}.$
They are conditionally independent given the agent’s hidden effort $a_{t}$:
$x_{t}\perp\widetilde{x}_{t}|a_{t}.$
The causal relationship between variables are summarized in Figure 1, with the dashed items included.
@@
Agents are sufficiently diverse in their talent. Fix any $\beta\in\mathcal{B}$, we have that the following holds
$\lambda_{\min}\mathbb{E}_{c(\cdot)\sim\mathcal{D}}[a(\beta;c(\cdot))^{\top}a(\beta;c(\cdot))]\geq\lambda_{0}.$
The first part of the assumption requires repeated observations, which is often realistic in practical scenarios. Some literature in economics also study the problem of repeated outputs (though under very different contexts, for more discussion see Chapter 10.2 in [Bolton and Dewatripont, 2004]).
The second part of the assumptions is essentially requireing that there is sufficient diverity in the agent’s ability to complete tasks. We give a concrete example satisfying assumption below.
Example 1.
Consider the cost functions of the agents have diagonal quadratic forms. Specifically, the cost function of the agent is parameterized by a vector $\kappa\in(\mathbb{R}^{d})^{+}$, so that the cost function of the agent is $c(a)=\frac{1}{2}a^{\top}\mathtt{diag}(\kappa)^{-1}a$. Then, the agent’s best response to a contract $\beta$ is the maximizer of $\langle\beta,a\rangle-c(a)$, which is equal to
$a=\kappa\odot\beta.$
We assume that the agent cost function in each round, parameterized by $\kappa_{t}$, is drawn from an unknown distribution. The distribution satisfies the following.
$\lambda_{\min}\mathbb{E}_{\kappa\sim\mathcal{D}}[\kappa\kappa^{\top}]\geq\lambda.$
Then we know that
$\lambda_{\min}\mathbb{E}[a(\beta)a^{\top}(\beta)]=c_{0}^{2}\cdot\lambda.$
Hence the diversity condition is satisfied. Note that this model has also appeared in the prior work [Thiele, 2010].
In addition to the above two assumptions, we require a relatively mild boundedness assumption as below.
Assumption 7.
Fix any $\beta\in B$. We have the following for any agent type: $\norm{a}_{2}^{2}\leq K_{0}$. Here $a$ is the best response to contract $\beta$.
7.1 Offline Setting: An Alternative GMM Estimator
The causal relationship is summarized in Figure 1 with the dashed items included. It can be observed that the second set of observations can serve as an instrumental variable for the other, and that we have: